# Three short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b and TOI-6086 b

### v1

@@
One of the main goals of the NASA’s TESS (Transiting Exoplanet Survey Satellite) mission is the discovery of Earth-like planets around nearby M-dwarf stars. Here, we present the discovery and validation of three new short-period Earth-sized planets orbiting nearby M-dwarfs: TOI-5720 b, TOI-6008 b and TOI-6086 b. We combined TESS data, ground-based multi-color light curves, ground-based optical and near-infrared spectroscopy, and Subaru/IRD RVs data to validate the planetary candidates and constrain the physical parameters of the systems. In addition, we used archival images, high-resolution imaging, and statistical validation techniques to support the planetary validation. TOI-5720 b is a planet with a radius of $R_{p}=1.09\pm 0.07\leavevmode\nobreak\ R_{\oplus}$ orbiting a nearby (23 pc) M2.5 host, with an orbital period of $P=1.4344555\pm 0.0000036$ days. It has an equilibrium temperature of $T_{\rm eq}=708\pm 19$ K (assuming a null albedo) and an incident flux of $S_{p}=41.7\pm 4.5\leavevmode\nobreak\ S_{\oplus}$. TOI-6008 b has a period of $P=0.8574347\pm 0.0000424$ day, a radius of $R_{p}=1.03\pm 0.05\leavevmode\nobreak\ R_{\oplus}$, an equilibrium temperature of $T_{\rm eq}=707\pm 19$ K and an incident flux of $S_{p}=41.5\pm 4.5\leavevmode\nobreak\ S_{\oplus}$. The host star (TOI-6008) is a nearby (36 pc) M5 with an effective temperature of $T_{\rm eff}=3075\pm 75\leavevmode\nobreak\ K$. Based on the RV measurements collected with Subaru/IRD, we set a $3\sigma$ upper limit of $M_{p}<4M_{\oplus}$, thus ruling out a star or brown dwarf as the transiting companion. TOI-6086 b orbits its nearby (31 pc) M3 host star ($T_{\rm eff}=3200\pm 75$ K) every $1.3888725\pm 0.0000827$ days, and has a radius of $R_{p}=1.18\pm 0.07\leavevmode\nobreak\ R_{\oplus}$, an equilibrium temperature of $T_{\rm eq}=634\pm 16$ K and an incident flux of $S_{p}=26.8\pm 2.7\leavevmode\nobreak\ S_{\oplus}$. Additional high precision radial velocity measurements are needed to derive the planetary masses and bulk densities, and to search for additional planets in the systems. Moreover, short-period earth-sized planets orbiting around nearby M-dwarfs are suitable targets for atmospheric characterization with the James Webb Space Telescope ( JWST) through transmission and emission spectroscopy, and phase curve photometry.
@@ 1 Introduction
M dwarfs are stars with effective temperatures $<3800$ K (e.g. Reid et al. (1995); Nutzman & Charbonneau (2008); Kaltenegger & Traub (2009); Winters et al. (2014)). They are the most common stars in our galaxy (Henry et al. 1994; Kirkpatrick et al. 1999). M-dwarf stars are excellent targets to search for transiting Earth-like planets. Such systems offer a rare opportunity for exploring their physical, orbital and atmospheric parameters, due to their low masses, small sizes and low luminosities. The combination of the high equilibrium temperature of the planet, low mass, small radius and infrared brightness of the host star, making the short-period planets around M dwarfs exciting targets for detailed atmospheric characterization using the emission and transmission spectroscopy, and phase curve photometry with JWST (Zieba et al. 2023; Greene et al. 2023). Those planets with especially short periods (of order one day) may provide us some additional information into the formation, evolution and interior structure of exoplanetary systems (Winn et al. 2018). Moreover, they are also excellent systems to derive the slope of the radius valley (see, e.g., Fulton et al. (2017)), and determine star-planet interactions, such as the tidal dissipation’ effect within their host stars (see, e.g., Aller et al. (2020)).
In this context, we present the discovery and validation of three short-period Earth-sized TESS planets around the nearby M dwarfs TOI-5720, TOI-6008 and TOI-6086. By combining TESS data, and ground-based photometric and spectroscopic measurements, we find that TOI-5720 b, TOI-6008 b and TOI-6086 b are $R_{p}=1.09\pm 0.07\leavevmode\nobreak\ R_{\oplus}$, $R_{p}=1.03\pm 0.05\leavevmode\nobreak\ R_{\oplus}$ and $R_{p}=1.18\pm 0.07\leavevmode\nobreak\ R_{\oplus}$ with orbital periods of $P=1.4344555\pm 0.0000036$, $P=0.8574347\pm 0.0000424$ and $P=1.3888725\pm 0.0000827$ days, respectively.
The paper is organized as follows. Section 2 describes the TESS and ground-based observations. Section 3 describes the stellar characterisation of the host-stars using spectral energy distributions (SEDs) and spectroscopic observations. Section 4 presents the validation of the transit signals in the light curves. The planet searches and detection limits from the TESS photometry are presented in Section 5. Section 6 presents our global analysis of all available photometric and spectroscopic datasets of the planetary systems, allowing us to derive the physical parameters of the systems. Our discussion and conclusion are presented in Section 7.
@@ 2.1 TESS  photometry
The host stars TIC 230055368 (TOI-5720), TIC 286201103 (TOI-6008) and TIC 18318288 (TOI-6086) were observed by the TESS  (Ricker et al. 2015) mission with 2-minutes cadence. TOI-5720 was observed in sectors 22 (2020 Feb 18 to 2020 Mar 18) and 49 (2022 Feb 26 to 2022 Mar 26). TOI-6008 was observed in sectors 15 (2019 Aug 15 to 2019 Sep 11), 16 (2019 Sep 11 to 2019 Oct 07), 41 (2021 Jul 23 to 2021 Aug 20), 55 (2022 Aug 05 to 2022 Sep 01), 56 (2022 Sep 01 to 2022 Sep 30) and 75 (2024 Jan 30 to 2024 Feb 26). TOI-6086 was observed in sectors 25 (2020 May 13 to 2020 Jun 08), 26 (2020 Jun 08 to 2020 Jul 04), 52 (2022 May 18 to 2022 Jun 13) and 53 (2022 Jun 13 to 2022 Jul 09). To perform TESS data modeling, we retrieved the Presearch Data Conditioning light curves (PDC-SAP, Stumpe et al. (2012); Smith et al. (2012); Stumpe et al. (2014) constructed by the TESS Science Processing Operations Center (SPOC; Jenkins et al. (2016)) at Ames Research Center from the Mikulski Archive for Space Telescopes. The PDC-SAP light curves have been calibrated for any instrument systematics and crowding effects. The TESS field-of-view for each target and photometric apertures used with the location of nearby Gaia DR3 sources (Gaia Collaboration et al. 2021a) are shown in Figure 1. Figures 2, 3 and 16 show the TESS photometric data for each target.
Refer to captionRefer to captionRefer to caption
Figure 1: TESS target pixel file images of TOI-5720 (left panel) TOI-6008 (middle panel) and TOI-6086 (right panel) observed in Sectors 22, 15 and 25, respectively. The mages are made by the tpfplotter package (Aller et al. 2020). Red dots and red shaded polygons show the position of Gaia DR3 sources and photometric apertures used to extract the photometric measurements, respectively.
Refer to captionRefer to captionRefer to caption
Figure 2: TESS PDC-SAP flux extracted from the 2-minutes cadence data of TOI-5720 (top panel), TOI-6008 (middle panel) and TOI-6086 (bottom panel). TOI-5720 was observed in Sectors 22 and 49, TOI-6008 was observed in Sectors 15, 16, 41, 55, 56 and 75, and TOI-6086 was observed in Sectors 25, 26, 52 and 53. Light grey points show the 2-minutes cadence data, and the black points shows the flux in 30-minutes bins. The transit locations of TOI-5720 b, TOI-6008 b and TOI-6086 b are shown with the blue triangles.
Refer to captionRefer to captionRefer to caption
@@ 2.2 Ground-based photometric follow-up
We performed ground-based observations follow-up of TOI-5720, TOI-6008 and TOI-6086 as part of the TESS Follow-up Observing Program (TFOP) in order to (i) rule out the NEBs (Nearby Eclipsing Binaries) as potential sources of the TESS detection, (ii) confirm the transit event on the target star, (iii) check the chromaticity of the transit depth measurement, (iv) refine the transit ephemerides. We used the TESS Transit Finder tool (Jensen 2013) in order to schedule the photometric follow-up observations. All ground-based photometric observations are summarized in Table 1.
@@
We used the LCOGT (Las Cumbres Observatory Global Telescope; Brown et al. (2013)) 2.0m Faulkes Telescope North located at Haleakala Observatory in Hawaii to observe TOI-5720 b, TOI-6008 b and TOI-6086 b. The telescope is equipped with the MuSCAT3 multi-band imager (Narita et al. 2020). One full transit of TOI-5720 b was observed on UTC May 12 2023, three full transits of TOI-6008 b were observed on UTC May 07, 13 and 31 2023, and two full transits of TOI-6086 b were observed on UTC July 06 and 31 2023. All observations were carried out simultaneously with the Sloan-$g^{\prime}$, -$r^{\prime}$, -$i^{\prime}$ and Pan-STARRS-$z_{\mathrm{s}}$ filters. The stellar fluxes were extracted using uncontaminated photometric apertures (see Table 1).
@@ 2.2.2 MuSCAT2
We observed one full transit of TOI-6008 b on UTC July 27 2023 and one full transit of TOI-6086 b on UTC July 29 2023 with the 1.52m Telescopio Carlos Sanchez (TCS) at the Teide Observatory in Spain. The telescope is equipped with the MuSCAT2 multicolor imager (Narita et al. 2019). Both of transits were observed simultaneously in the Sloan-$g$, -$r$, -$i$, and Pan-STARRS-$z_{\mathrm{s}}$. The stellar fluxes were extracted using an uncontaminated photometric apertures (see Table 1). The data calibration and aperture photometry were performed using the MuSCAT2 photometry pipeline (Parviainen et al. 2020).
@@
Two full transits of TOI-6086 b were observed with LCOGT-1m0 network on UTC March 10 and 20 2023 in the Sloan-$i^{\prime}$ filter. The telescopes are equipped with $4096\times 4096$ SINISTRO Cameras with a pixel scale of $0.389\arcsec$ per pixel and a FOV of $26^{\prime}\times 26^{\prime}$. The science images were calibrated by the standard LCOGT BANZAI pipeline and aperture photometry was performed using AstroImageJ software. Both of observations are used for checking nearby eclipsing binaries, but they are not included in our global analysis because of the low signal-to-noise ratio.
@@
Table 1: Table shows the observational parameters for TOI-5720 b, TOI-6008 b and TOI-6086 b: date of observation, filter used, telescope, exposure time(s), and FWHM of the point-spread function and photometric aperture. Analysis parameters: the selected baseline model, red noise $\beta_{r}$, white noise $\beta_{w}$ and $CF$ (Correction Factor). The baseline model is shown as a polynomial $p()$ of systematics parameters, with $t=$ time, $x=$ and $y=$ positions of the star on the detector, $a=$ airmass, $b=$ background, and $f=$ full-with at half maximum of the PSF.
@@ 2.2.5 IRTF/SpeX Spectroscopy
We used the SpeX spectrograph (Rayner et al. 2003) on the 3.2m-NASA IRTF (Infrared Telescope Facility) to collect medium-resolution near-infrared spectra of TOI-6086 on UT 2023 May 04 and of TOI-5720 and TOI-6008 on UT 2023 June 05. Thin cirrus clouds were present on both nights. Seeing was ${\prime\prime}$0 on the May night and ${\prime\prime}$6 on the June night. On both nights, we used the short-wavelength cross-dispersed (SXD) mode with the $0\aas@@fstack{\prime\prime}3\times 15^{\prime\prime}$ slit aligned to the parallactic angle, providing spectra covering 0.80–2.42 $\mu$m with a resolving power of $R{\sim}2000$ and an average of 2.5 pixels per resolution element. Nodding in an ABBA pattern, we collected six integrations of TOI-5720 and TOI-6008 and eight integrations of TOI-6086 with integration times of 70.0 s, 70.0 s, and 149.7 s, respectively. After each science observation, we collected a set of standard SXD flat-field and arc-lamp exposures and six integrations of A0 V standard stars. We reduced the data with Spextool v4.1 (Cushing et al. 2004), following the standard approach outlined in the Spextool User’s Manual [^3] Available at http://irtfweb.ifa.hawaii.edu/~spex/observer/.. The final spectra have median signal-to-noise ratios per pixel of 124, 109, and 133 for TOI-5720, TOI-6008, and TOI-6086, respectively (see Figure 5).
Figure 5: SpeX SXD spectra of TOI-5720 (top), TOI-6008 (middle), and TOI-6086 (bottom). The target spectra (blue) are shown alongside the spectrum of the best-fit standard star (grey). Regions of high telluric absorption are shaded, and strong spectral features of M dwarfs are highlighted.

### v2

# Three short-period Earth-sized planets around M dwarfs discovered by TESS: TOI-5720 b, TOI-6008 b, and TOI-6086 b

@@
One of the main goals of the NASA Transiting Exoplanet Survey Satellite (TESS) mission is the discovery of Earth-like planets around nearby M-dwarf stars. We present the discovery and validation of three new short-period Earth-sized planets orbiting nearby M dwarfs: TOI-5720 b, TOI-6008 b, and TOI-6086 b. We combined TESS data, ground-based multicolor light curves, ground-based optical and near-infrared spectroscopy, and Subaru/IRD radial velocity data to validate the planetary candidates and constrain the physical parameters of the systems. In addition, we used archival images, high-resolution imaging, and statistical validation techniques to support the planetary validation. TOI-5720 b is an Earth-sized planet with a radius of $R_{p}=1.09\pm 0.07\leavevmode\nobreak\ R_{\oplus}$. It orbits a nearby (36 pc) M2.5 host with an orbital period of $P=1.4344555\pm 0.0000036$ days. It has an equilibrium temperature of $T_{\rm eq}=708\pm 19$ K (assuming a null albedo) and an incident flux of $S_{p}=41.7\pm 4.5\leavevmode\nobreak\ S_{\oplus}$. TOI-6008 b is a short-period planet of $P=0.8574347\pm 0.0000424$ day. It has a radius of $R_{p}=1.03\pm 0.05\leavevmode\nobreak\ R_{\oplus}$, an equilibrium temperature of $T_{\rm eq}=707\pm 19$ K, and an incident flux of $S_{p}=41.5\pm 4.5\leavevmode\nobreak\ S_{\oplus}$. The host star (TOI-6008) is a nearby (23 pc) M5 with an effective temperature of $T_{\rm eff}=3075\pm 75\leavevmode\nobreak\ K$. Based on the radial velocity measurements collected with Subaru/IRD, we set a $3\sigma$ upper limit of $M_{p}<4M_{\oplus}$, thus ruling out a star or brown dwarf as the transiting companion. TOI-6086 b orbits its nearby (32 pc) M3 host star ($T_{\rm eff}=3200\pm 75$ K) every $1.3888725\pm 0.0000827$ days and has a radius of $R_{p}=1.18\pm 0.07\leavevmode\nobreak\ R_{\oplus}$, an equilibrium temperature of $T_{\rm eq}=634\pm 16$ K, and an incident flux of $S_{p}=26.8\pm 2.7\leavevmode\nobreak\ S_{\oplus}$. Additional high-precision radial velocity measurements are needed to derive the planetary masses and bulk densities and to search for additional planets in the systems. Moreover, short-period Earth-sized planets orbiting around nearby M dwarfs are suitable targets for an atmospheric characterization with the James Webb Space Telescope through transmission and emission spectroscopy and phase-curve photometry.
@@ 1 Introduction
M dwarfs are stars with effective temperatures $<3800$ K (e.g. Reid et al. (1995); Nutzman & Charbonneau (2008); Kaltenegger & Traub (2009); Winters et al. (2014)). They are the most common stars in our Galaxy (Henry et al. 1994; Kirkpatrick et al. 1999). M-dwarf stars are excellent targets for a search for transiting Earth-like planets. These systems offer a rare opportunity for exploring their physical, orbital, and atmospheric parameters due to their low masses, small sizes and low luminosities. The combination of the high-equilibrium temperature of the planet, low mass, small radius, and infrared brightness of the host star make the short-period planets around M dwarfs exciting targets for a detailed atmospheric characterization using emission and transmission spectroscopy and phase-curve photometry (Zieba et al. 2023; Greene et al. 2023). These planets with especially short periods (of about one day) may provide some additional information about the formation, evolution, and interior structure of exoplanetary systems (Winn et al. 2018). Moreover, they are also excellent systems based on which the slope of the radius valley can be derived (see, e.g., Fulton et al. (2017)), and star-planet interactions, such as the tidal dissipation’ effect within their host stars (see, e.g., Aller et al. (2020)) can be determined.
In this context, we present the discovery and validation of three short-period Earth-sized TESS planets around the nearby M dwarfs TOI-5720, TOI-6008, and TOI-6086. By combining TESS data, and ground-based photometric and spectroscopic measurements, we find that TOI-5720 b, TOI-6008 b and TOI-6086 b have radii of $R_{p}=1.09\pm 0.07\leavevmode\nobreak\ R_{\oplus}$, $R_{p}=1.03\pm 0.05\leavevmode\nobreak\ R_{\oplus}$ and $R_{p}=1.18\pm 0.07\leavevmode\nobreak\ R_{\oplus}$ with orbital periods of $P=1.4344555\pm 0.0000036$, $P=0.8574347\pm 0.0000424$, and $P=1.3888725\pm 0.0000827$ days, respectively.
The paper is organized as follows. Section 2 describes the TESS and ground-based observations. Section 3 describes the stellar characterization of the host stars using spectral energy distributions (SEDs) and spectroscopic observations. Section 4 presents the validation of the transit signals in the light curves. The planet searches and detection limits from the TESS photometry are presented in Section 5. Section 6 presents our global analysis of all available photometric and spectroscopic datasets of the planetary systems, allowing us to derive the physical parameters of the systems. Our discussion and conclusion are presented in Section 7.
@@ 2.1 TESS  photometry
The host stars TIC 230055368 (TOI-5720), TIC 286201103 (TOI-6008) and TIC 18318288 (TOI-6086) were observed by the TESS (Ricker et al. 2015) mission with a cadence of 2-minutes. TOI-5720 was observed in sectors 22 (2020 February 18 to 2020 March 18) and 49 (2022 February 26 to 2022 March 26). TOI-6008 was observed in sectors 15 (2019 August 15 to 2019 September 11), 16 (2019 September 11 to 2019 October 07), 41 (2021 July 23 to 2021 August 20), 55 (2022 August 05 to 2022 September 01), 56 (2022 September 01 to 2022 September 30) and 75 (2024 January 30 to 2024 February 26). TOI-6086 was observed in sectors 25 (2020 May 13 to 2020 June 08), 26 (2020 June 08 to 2020 July 04), 52 (2022 May 18 to 2022 June 13) and 53 (2022 June 13 to 2022 July 09). To model the TESS data, we retrieved the presearch data conditioning light curves (PDC-SAP; Stumpe et al. (2012); Smith et al. (2012); Stumpe et al. (2014)) constructed by the TESS Science Processing Operations Center (SPOC; Jenkins et al. (2016)) at the Ames Research Center from the Mikulski Archive for Space Telescopes. The PDC-SAP light curves were calibrated for any instrument systematics and crowding effects. The TESS field of view for each target and photometric apertures used with the location of nearby Gaia DR3 sources (Gaia Collaboration et al. 2021a) are shown in Figure 1. Figures 2, 3 and 15 show the TESS photometric data for each target.
Figure 1: TESS target pixel file images of TOI-5720 (left panel) TOI-6008 (middle panel) and TOI-6086 (right panel) observed in sectors 22, 15 and 25, respectively. The images are made by the tpfplotter package (Aller et al. 2020). Red dots and red shaded polygons show the positions of Gaia DR3 sources and photometric apertures used to extract the photometric measurements, respectively.
Figure 2: TESS PDC-SAP flux extracted from the 2-minute cadence data of TOI-5720 (top panel), TOI-6008 (middle panel), and TOI-6086 (bottom panel). TOI-5720 was observed in sectors 22 and 49, TOI-6008 was observed in sectors 15, 16, 41, 55, 56 and 75, and TOI-6086 was observed in sectors 25, 26, 52 and 53. The light grey points show the 2-minute cadence data, and the black points shows the flux in 30-minutes bins. The transit locations of TOI-5720 b, TOI-6008 b, and TOI-6086 b are shown with the blue triangles.
@@ 2.2 Ground-based photometric follow-up
We performed ground-based observation follow-up of TOI-5720, TOI-6008 and TOI-6086 as part of the TESS follow-up observing program (TFOP) in order to (i) rule out the nearby eclipsing binaries (NEBs) as potential sources of the TESS detection, (ii) confirm the transit event on the target star, (iii) check the chromaticity of the transit depth measurement, and (iv) refine the transit ephemerides. We used the TESS transit finder tool (Jensen 2013) in order to schedule the photometric follow-up observations. All ground-based photometric observations are summarized in Table 1.
@@
We used the Las Cumbres Observatory Global Telescope (LCOGT; Brown et al. (2013)) 2.0m Faulkes Telescope North located at Haleakala Observatory in Hawaii to observe TOI-5720 b, TOI-6008 b and TOI-6086 b. The telescope is equipped with the MuSCAT3 multiband imager (Narita et al. 2020). One full transit of TOI-5720 b was observed on UTC May 12 2023, three full transits of TOI-6008 b were observed on UTC May 7, 13 and 31 2023, and two full transits of TOI-6086 b were observed on UTC July 6 and 31 2023. All observations were carried out simultaneously with the Sloan-$g^{\prime}$, -$r^{\prime}$, -$i^{\prime}$ and Pan-STARRS-$z_{\mathrm{s}}$ filters. The stellar fluxes were extracted using uncontaminated photometric apertures (see Table 1).
@@ 2.2.2 MuSCAT2
We observed one full transit of TOI-6008 b on UTC July 27 2023 and one full transit of TOI-6086 b on UTC July 29 2023 with the 1.52m Telescopio Carlos Sanchez (TCS) at the Teide Observatory in Spain. The telescope is equipped with the MuSCAT2 multicolor imager (Narita et al. 2019). Both transits were observed simultaneously in the Sloan-$g$, -$r$, -$i$, and Pan-STARRS-$z_{\mathrm{s}}$. The stellar fluxes were extracted using an uncontaminated photometric aperture (see Table 1). The data calibration and aperture photometry were performed using the MuSCAT2 photometry pipeline (Parviainen et al. 2020).
@@
Two full transits of TOI-6086 b were observed with LCOGT-1m0 network on UTC March 10 and 20 2023 in the Sloan-$i^{\prime}$ filter. The telescopes are equipped with $4096\times 4096$ SINISTRO cameras with a pixel scale of $0.389\arcsec$ per pixel and an FOV of $26^{\prime}\times 26^{\prime}$. The science images were calibrated by the standard LCOGT BANZAI pipeline, and aperture photometry was performed using the software AstroImageJ. Both observations are used to check nearby eclipsing binaries, but they were not included in our global analysis because of the low signal-to-noise ratio (S/N).
@@
Table 1: Observational parameters for TOI-5720 b, TOI-6008 b, and TOI-6086 b: Date of the observation, filter, telescope, exposure time(s), and FWHM of the point-spread function and photometric aperture. The analysis parameters are the selected baseline model, red noise $\beta_{r}$, white noise $\beta_{w}$ and $CF$ (Correction Factor). The baseline model is shown as a polynomial $p()$ of systematic parameters, with $t$ is the time, $x$ and $y$ are the positions of the star on the detector, $a$ is the airmass, $b$ is the background, and $f$ is the full with at half maximum of the PSF.
@@ 2.2.5 IRTF/SpeX spectroscopy
We used the SpeX spectrograph (Rayner et al. 2003) on the 3.2m-NASA Infrared Telescope Facility (IRTF) to collect medium-resolution near-infrared spectra of TOI-6086 on UT 2023 May 4 and of TOI-5720 and TOI-6008 on UT 2023 June 5. Thin cirrus clouds were present on both nights. Seeing was ${\prime\prime}$0 on the May night and ${\prime\prime}$6 on the June night. On both nights, we used the short-wavelength cross-dispersed (SXD) mode with the $0\aas@@fstack{\prime\prime}3\times 15^{\prime\prime}$ slit aligned to the parallactic angle, providing spectra covering 0.80–2.42 $\mu$m with a resolving power of $R{\sim}2000$ and an average of 2.5 pixels per resolution element. Nodding in an ABBA pattern, we collected six integrations of TOI-5720 and TOI-6008 and eight integrations of TOI-6086 with integration times of 70.0 s, 70.0 s, and 149.7 s, respectively. After each science observation, we collected a set of standard SXD flat-field and arc-lamp exposures and six integrations of A0 V standard stars. We reduced the data with Spextool v4.1 (Cushing et al. 2004), following the standard approach outlined in the Spextool User Manual [^3] Available at http://irtfweb.ifa.hawaii.edu/~spex/observer/.. The final spectra have median signal-to-noise ratios per pixel of 124, 109, and 133 for TOI-5720, TOI-6008, and TOI-6086, respectively (see Figure 5).
Figure 5: SpeX SXD spectra of TOI-5720 (top), TOI-6008 (middle), and TOI-6086 (bottom). The target spectra (blue) are shown alongside the spectrum of the best-fit standard star (gray). Regions of high telluric absorption are shaded, and strong spectral features of M dwarfs are highlighted.